What is the acceleration from 5.0 m/s to 15 m/s south in 4.0 second
1
Expert's answer
2020-10-21T10:15:48-0400

Acceleration is defined as change in velocity per unit time.

In the given question, initial velocity is 5 m/s

Final velocity is 15 m/s

Time taken is 4 s.

Then acceleration is given by,

"a=\frac{v-u}{t}=\frac{15-5}{4}=2.5m/s^2" in South direction.